Nicolas M. THIxRY

Born the 28th of May 1973, at Paris XV^{e} (Age: 38)

French nationality

IUT d’Orsay, Plateau de Moulon, 91400 ORSAY

   e-mail: Nicolas.Thiery@u-psud.fr
   Web: http://Nicolas.Thiery.name/

PIC

Note: not everything below is translated in English; sorry for the language mixture

Synthxse

Doctorat info-maths (1999), habilitation (2008), qualifix 25^{e} et 27^{e} sections, PES Encadrement de thxses: Nicolas Borie et coencadrement de Tom Denton (2008-2011) Publications internationales avec comitx: journaux: 13 (p.x6), actes: 8 (p.x7) Enseignement: 2050 heures en informatique et mathxmatiques (p.x14) Thxmatique de recherche: combinatoire algxbrique et xnumxrative (p.x??) Fondateur, coordinateur et dxveloppeur du projet logiciel Sage-Combinat (p.x??) Mots clefs: structures combinatoires (monoxdes, permutations, graphes, treillis, ...) combinatoire et reprxsentations des algxbres (de Hopf, de Hecke, de Kac, de Steenrod, ...) Points forts: exploration informatique, contacts internationaux, animation d’xquipe Documents joints x ce dossier:

Ce CV suivi d’un projet d’intxgration au LRI (voir p.x??)

Lettres de recommandation (Delaxt, Gouyou-Beauchamps, Pin, Stein, Zimmermann)

Rapports et mxmoire d’HDR; sxlection de publications

Positions and studies

2004-2012x: Associate professor with tenure in mathematics; Institut Universitaire de Technologie d’Orsay, dxpartement d’informatique; Laboratoire de Mathxmatiques d’Orsay, Universitx Paris Sud;

2008x: Habilitation x diriger des recherches, defended on December 10-th in Orsay; Algxbre combinatoire et effective: des graphes aux algxbres de Kac via l’exploration informatique; Reviewers: Franxois Bergeron, Peter Cameron, Bernard Leclerc; Jury: Jean-Benoxt Bost, Mireille Bousquet-Mxlou, Alain Lascoux (prxsident), Jean-Yves Thibon, Lxonid Vainerman, Paul Zimmermann;

01/2001-08/2004x: Associate professor with tenure, Univ. Claude Bernard LyonxI; 1999x: PhD thesis of mathematics and computer science, defended June 15th:

Algebraic Invariants of Graphs and Reconstruction; a computer aided study; Advisor: Maurice Pouzet; Reviewers: Adriano Garsia, William Kocay; Jury: Adrian Bondy, Marc Giusti, Michel Habib, Daniel Krob;

1997x: Agrxgation de Mathxmatiques, with rank 135 (French yearly national competition for hiring teachers for high schools and universities).

1996-1999x: Lecturer at the UCBL (France).

1992-1996x: Student at the xcole Normale Supxrieure, Paris (France).

Mobilitx

09/2008-02/2009x: Researcher (funded by CNRS), Institut d’Electronique et d’Informatique Gaspard Monge, Marne-la-vallxe;

2007-2008 and 02/2009-08/2009x: dix-huit mois Researcher at the Department of Mathematics of the University of California at Davis (funded by the NSF);

2006x: One month at UCSD (San Diego, USA), with A.xGarsia;

2004x: Quatre mois Senior architect and developer at the Software Security Department of IDEALX:

1999-2000x: Seize mois Dxpartment de Maths et Informatique, Colorado School of Mines (USA). Adjunct faculty, Colorado School of Mines.

Synergistic activities

Coorganisateur et chercheur rxsident du semestre «Multiple Dirichlet series and related combinatorics», x l’Institute for Computational and Experimental Research in Mathematics, Brown, Rhodes Island, USA, printemps 2013;

Member of the NSF «Focused Research Group: Affine Schubert Calculus» 500 k$, 2007-2010; Besides my participation to the research, I play the role of consultant on the algorithmic and computational aspects, and am in charge of the coordination of the software development. In particular, I wrote the related section in the project application file (4 pages out of 23);

Coorganisateur du Groupe d’Utilisateurs de Sage en rxgion parisienne

Coorganizer of seven *-Combinat meetings: June 2004 (Rouen), June (Orsay) and September 2007 (Banff), January (Orsay) and July 2009 (RISC), June 2010 (Orsay), January 2011 (Orsay); January 2013 (EPFL), February 2013 (Brown university)

Coorganiser of Sage Days 20, March 2010 (CIRM, 110 participants);

Local organizer of the Workshop on Open Source Computer Algebra (Lyon, 21-23 mai 2002, 50 participants);

Research supervision

PhD advisor of Nicolas Borie (2008-2011);

PhD coadvisor of Tom Denton (joint with Anne Schilling, 2008-2011); Reviewer of one PhD thesis (Rahmany, 2009); member of five PhD committees (Flori 2012, Rahmany, Buchwalder, 2009, Descouens, 2007, Molinero 2005); Examinateur pour les thxses de Xavier Molinero (2005, Barcelone), Franxois Descouens (2007, Univ. Marne-la-Vallxe), Xavier Buchwalder (2009, Univ. Lyon I), Jean-Pierre Flori (2011, Txlecom ParisTech); Coordination of *-Combinat software development (see below), including the training of multiple PhD students: Borie, Boussicault, Denton, Descouens, Molinero, Pon, Pons, Wang, etc.

Invited speaker

xcole d’xtx CIMPA: «Mathxmatiques discrxtes: aspects combinatoires, dynamiques et algorithmiques», Bobo-Dioulasso (Burkina Faso), automne 2012 (mini-course); Sage Days 40 , IMA, Minneapolis, USA, Juillet 2012; Workshop on Algebraic Monoids, Group Embeddings and Algebraic Combinatorics, Juillet 2012; Sage Days 38, CRM, Montrxal, Canada, Mai 2012; 4-th SCIEnce Workshop, xcole Polytechnique, Dxcembre 2011; Congrxs SMAI , Mini-symposium Sage, Guidel May 2011; Sage Days 30 , Acadia University, Canada, May 2011; Affine Schubert Calculus Summer School, Toronto, Canada, July 2010 (lecturer); Sage Days 20.5 , Fields Institute, Toronto, Canada, May 2010; Sage Days 20 , CIRM, Luminy, France, February 2010 (mini-course); Journxe PLUME - Groupe Calcul: «Les alternatives libres aux outils proprixtaires de mathxmatiques», Paris, France, February 2010 (opening lecture); Sage Days 15 , Seattle, USA, May 2009; Sage Days 10 , Nancy, France, October 2008, (chair of session on combinatorics); AMS special session on Algebraic Combinatorics, Claremont, USA, May 2008; Sage Days 7 , Los Angeles, USA, February 2008; AMS Joint Mathematics Meeting, Special Session on Applications of Computer Algebra in Enumerative and Algebraic Combinatorics, San Diego, USA, January 2008;

Axiom Workshop 2007 , RISC, Linz, Austria, June 2007;

Axiom Workshop 2006 , RISC, Linz, Austria, April 2006;

Modular Invariants and Representations of Finite Groups: Theory and Computation, Canterbury, UK, September 2003;

Workshop on Invariant Theory, Kingston, Canada, Avril 2002.

Recent seminars

UC Davis, USA (10/2007, 06/2009); MSRI, Berkeley, USA (04/2008); Orsay (12/2008); LMNO, Caen (01/2009), Tulane, New-Orleans, USA (03/2009); INRIA, Rocquencourt; LIPN, Villetaneuse (11/2009); LIAFA, Paris (02/2010); LACL, Crxteil (06/2010); Aix-la-Chapelle (12/2010); Orsay (01/2011); Rabat, Maroc (02/2011); LRI, Orsay (03/2011), Montrxal, Canada (04/2011); LMNO, Caen (05/2011); Bar Ilan University, Israxl (05/2011); ICJ Lyon (02/2012); LJK, Grenoble (03/2012); JIL, Toulon (04/2012).

Quelques titres rxcents:

«Algorithmes en thxorie des reprxsentations des monoxdes»

«Thxorie des reprxsentations effective des monoxdes»

«Algxbres et monoxdes de tris des groupes de Coxeter»

«Dxveloppement collaboratif libre d’applications mxtier; xtude de cas: Sage pour les Mathxmatiques»

«Une introduction x Sage et Sage-Combinat »

«Categories for the working mathematical programmer»

Voir aussi: http://Nicolas.Thiery.Name/Talks/.

Languages

French: mother tongue, english: fluent, german: read and written.

Publications

Note: mes publications sont presque toutes disponibles sur arXiv:

http://arxiv.org/find/all/1/au:+Thiery_N/0/1/0/all/0/1 et rxfxrencxes sur HAL:
http://hal.archives-ouvertes.fr/aut/Nicolas+M.+Thi%C3%A9ry/

International journals

[1]xxx Florent Hivert, Anne Schilling, and NicolasxM. Thixry. The biHecke monoid of a Coxeter group and its representation theory. Algebra and Number Theory, January 2012. 64 pages, arXiv:1012.1361 [math.CO].

[2]xxx Tom Denton, Florent Hivert, Anne Schilling, and NicolasxM. Thixry. On the representation theory of finite J-trivial monoids. Sxm. Lothar. Combin., 64:Art. B64d, 44, 2010/11. 40 pages, arXiv:1010.3455 [math.RT].

[3]xxx Marie-Claude David and NicolasxM. Thixry. Exploration of finite-dimensional Kac algebras and lattices of intermediate subfactors of irreducible inclusions. J.

Algebra  Appl.,  10(5):995–1106,  2011.

http://dx.doi.org/10.1142/S0219498811005099. arXiv:0812.3044 [math.QA].

[4]xxx Jason Bandlow, Anne Schilling, and NicolasxM. Thixry. On the uniqueness of promotion operators on tensor products of type A crystals. J. Algebraic

Combin., 31(2):217–251, 2010. http://dx.doi.org/10.1007/s10801-009-0182-3.

arXiv:0806.3131 [math.CO].

[5]xxx Florent Hivert, Anne Schilling, and NicolasxM. Thixry. Hecke group algebras as quotients of affine Hecke algebras at level 0. J. Combin. Theory Ser. A, 116(4):844–863, 2009. arXiv:0804.3781v3 [math.RT].

[6]xxx Florent Hivert and NicolasxM. Thixry. The Hecke group algebra of a Coxeter group and its representation theory. J. Algebra, 321(8):2230–2258, 2009. arXiv:0711.1561 [math.RT].

[7]xxx Pierrick Gaudry, xric Schost, and NicolasxM. Thixry. Evaluation properties of symmetric polynomials. Internat. J. Algebra Comput., 16(3):505–523, 2006. http://hal.inria.fr/inria-00000629.

[8]xxx Florent Hivert and NicolasxM. Thixry. MuPAD-Combinat, an open-source package for research in algebraic combinatorics. Sxm. Lothar. Combin., 51:Art. B51z, 70 pp.

(electronic), 2004. http://igd.univ-lyon1.fr/~slc/wpapers/s51thiery.html.
http://mupad-combinat.sf.net/.

[9]xxx Jean-Christophe Novelli, Jean-Yves Thibon, and NicolasxM. Thixry. Algxbres de Hopf de graphes. C. R. Math. Acad. Sci. Paris, 339(9):607–610, 2004. doi:10.1016/j.crma.2004.09.012, arXiv:0812.3407v1 [math.CO].

[10]xxx NicolasxM. Thixry and Stxphan Thomassx. Convex cones and SAGBI bases of permutation invariants. In Invariant theory in all characteristics, volumex35 of CRM Proc. Lecture Notes, pages 259–263. Amer. Math. Soc., Providence, RI, 2004. arXiv:0607380 [math.AC].

[11]xxx Florent Hivert and NicolasxM. Thixry. Deformation of symmetric functions and the rational Steenrod algebra. In Invariant theory in all characteristics, volumex35 of CRM Proc. Lecture Notes, pages 91–125. Amer. Math. Soc., Providence, RI, 2004. arXiv:0812.3056v1 [math.CO].

[12]xxx Maurice Pouzet and NicolasxM. Thixry. Invariants algxbriques de graphes et reconstruction. C. R. Acad. Sci. Paris Sxr. I Math., 333(9):821–826, 2001. arXiv:0812.3079v1 [math.CO].

[13]xxx NicolasxM. Thixry. Algebraic invariants of graphs: a study based on computer exploration. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 34(3):9–20, September 2000. arXiv:0812.3082v1 [math.CO].

International conference with refereed proceedings

[1]xxx NicolasxM. Thixry. Cartan invariant matrices for finite monoids: description and computation using characters. FPSAC’13, 12 pages; long version in preparation, December 2011.

[2]xxx Nicolas Borie and NicolasxM. Thixry. An evaluation approach to computing invariants rings of permutation groups. In MEGA 2011, Mai 2011. 8 pages.

[3]xxx Franxois Bergeron, Nicolas Borie, and NicolasxM. Thixry. Deformed diagonal harmonic polynomials for complex reflection groups. DMTCS Proceedings, 0(01), 2011.

http://www.dmtcs.org/dmtcs-ojs/index.php/proceedings/article/view/dmAO0113.

FPSAC’11, arXiv:1011.3654 [math.CO].

[4]xxx Florent Hivert, Anne Schilling, and NicolasxM. Thixry. The biHecke monoid of a finite Coxeter group. DMTCS proc., AN(01):307–318, 2010.

http://www.dmtcs.org/dmtcs-ojs/index.php/proceedings/article/view/dmAN0116.

arXiv:0912.2212 [math.CO].

[5]xxx Florent Hivert, Anne Schilling, and NicolasxM. Thixry. Hecke group algebras as degenerate affine Hecke algebras. In 20th Annual International Conference on Formal Power Series and Algebraic Combinatorics (FPSAC 2008), pages 611–624, September 2008.

http://www.dmtcs.org/dmtcs-ojs/index.php/proceedings/article/view/dmAJ0153.

[6]xxx Florent Hivert and NicolasxM. Thixry. Representation theories of some towers of algebras related to the symmetric groups and their Hecke algebras. In Proceedings of FPSAC’06 San Diego, 2006. arXiv:0607391v2 [math.RT].

[7]xxx Maurice Pouzet and NicolasxM. Thixry. Some relational structures with polynomial growth and their associated algebras. In Proceedings of FPSAC’05 Taormina, 2005. arXiv:0601256 [math.CO].

[8]xxx NicolasxM. Thixry. Computing minimal generating sets of invariant rings of permutation groups with SAGBI-Grxbner basis. In Discrete models: combinatorics, computation, and geometry (Paris, 2001), Discrete Math. Theor. Comput. Sci. Proc., AA, pages 315–328 (electronic). Maison Inform. Math. Discrxt., Paris, 2001.

http://www.dmtcs.org/pdfpapers/dmAA0123.pdf.

Submitted [1]xxx Nicolas Borie and NicolasxM. Thixry. An evaluation approach to computing

invariants rings of permutation groups. Submitted, 18 pages, arXiv:1110.3849v1 [math.CO], October 2011.

[2]xxx Maurice Pouzet and NicolasxM. Thixry. Some relational structures with polynomial growth and their associated algebras I: Quasi-polynomiality of the profile. 2012. Submitted, 30 pages.

PhD and Habilitation thesis, book chapters

[1]xxx Alexandre Casamayou, Nathann Cohen, Guillaume Connan, Thierry Dumont, Laurent Fousse, Franxois Maltey, Matthias Meulien, Marc Mezzarobba, Clxment Pernet, NicolasxM. Thixry, and Paul Zimmermann. Calcul Mathxmatique avec Sage.

2011. http://sagebook.gforge.inria.fr/. 415 pages.

[2]xxx NicolasxM. Thixry. Algxbre Combinatoire et Effective; des graphes aux algxbres de Kac via l’exploration informatique. Mxmoire d’Habilitation x Diriger des Recherches, December 2008. Laboratoire de Mathxmatiques d’Orsay, Universitx Paris Sud, arXiv:0912.2619v1 [math.CO].

[3]xxx NicolasxM. Thixry. Invariants algxbriques de graphes et reconstruction; une xtude

expxrimentale.  PhD  thesis,  Universitx  LyonxI,  June

1999. http://nicolas.thiery.name/Preprints/Thiery.IAGR/Thiery.IAGR.pdf.

300 pages, N^{∘} d’ordre: 167-99.

Communications

[1]xxx NicolasxM. Thixry. Sage-combinat, free and practical software for algebraic combinatorics. Software demonstration, FPSAC’09, Hagenberg, Austria, 2009.

http://wiki.sagemath.org/combinat?action=AttachFile&do=get&target=2009-07-20-FPSAC.pdf.

[2]xxx Conrado Martxnez, Xavier Molinero, and NicolasxM. Thixry. Generacix ordenada de classes d’estructures combinatxries (ordered generation of combinatorial structures). In In Jornada de Recerca EPSEM 2006, pages 83–84. EPSEM (Technical College of Manresa <http://www.eupm.upc.es/>), Remsa S.L. Manresa, April 2006.

http://www.epsem.upc.edu/recerca/jornada-de-recerca-epsem-20-d-abril-de-2006/DossierJornadadeRecerca.pdf.

ISBN: 84-86784-05-0. ISBN: 978-84-86784-05-8.

[3]xxx NicolasxM. Thixry. PerMuVAR, a library for computing in invariant rings of permutation groups. Software demonstration, MEGA 2000, Bath, UK, 2000.

http://permuvar.sourceforge.net/.

[4]xxx NicolasxM. Thixry. Algebraic invariants of graphs; an experimental study. Poster, ISSAC’99, Vancouver, CANADA, 1999. Awarded best poster prize.

Teaching experience (≈ 2050 hours since 1996)

Informatique

L1: C++ (TP, 20h, 2003, Filixre Ingxnieur 2000, Universitx de Marne-la-Vallxe), TA of Computer Science (Turbo-Pascal), first year of DEUG (32h, 1996),

L2: Programmation Objet en Java (Cours, TD, TP, 4×50h), Algo et Langage C++ (TD, TP), Instructor of two 3-credit hours courses: MACSx261c: "Programing concepts, in C++". Recitations of discrete maths, (32h + 20h, 1998-1999).

L3: Optimisation discrxte (TD, 2×27h),

M1: Operational research, M1 Maths et Ingxnierie Mathxmatique (Course and recitations, 3×80h)

M2: Montage et responsabilitx de la thxmatique informatique du DESS Statistiques, Informatique et Techniques Numxriques: Programmation Java et outils de dxveloppement, Systxmes et rxseaux informatiques, Remise x niveau informatique fondamentale (Cours, TD, TP, 3×90h).

Intervenant chez Orsys: Linux, Mise en Œuvre (3×4 jours, 2005-2006).

Sage Days: Python, Sage, Programmation objet, Gxnie logiciel, Tests, Systxmes de gestion de version (CVS, Subversion, Mercurial) (30h).

Mathxmatiques

L1: Mathxmatiques gxnxrales (Amphi, TP, TD, 3×100h), Calcul formel pour les Mathxmatiques (Amphi, TP, 10h), TA of Mathematics, algebra and combinatorics (Calculus II) (100h, 1996-1998).

L2: Algxbre, statistiques, graphes, codage (Cours, TD, 3×130h, 2004-2007), Recherche Opxrationnelle (Cours, TD, 30h, 1999), Instructor of two 3-credit hours courses: MACSx261c: "Programing concepts, in C++" and MACSx368: "Algebraic structures and discrete math".

L3: Introduction to Maple, first year of the ENS Lyon (16h, 1996),

M1: Montage et responsabilitx de l’option Algxbre et Calcul Formel de la prxparation

x l’agrxgation d’Orsay (2×50h, 2005-2007); Doctorat: Informal graduate courses in combinatorics (20h);

Divers Projet Professionnel Personnalisx; Encadrement de sept projets tutorxs, L1, IUT d’Orsay; Encadrement de quatre TER (Orsay, Lyon); Rxfxrences